# Automatic Fact-Checking with Frame-Semantics

### v1

@@ 4.2 OECD-Related Statistics
For the OECD Case Study, we collect all of the data tables available on the OECD Data Explorer.  [^7] https://data-explorer.oecd.org/ We construct a SQLite database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totalling 4.1 billion cells.
@@
To identify relevant tables for data extraction, we apply semantic similarity models to select the five most similar tables based on the extracted frame elements (FEs). The query consists of one or more FEs derived from frame-semantic parsing, while the target is a specific table. We encode each table’s name and description as text for this purpose. Given the brevity of our queries and the length of the table descriptions, we utilize asymmetric semantic similarity models to ensure precise matches. To minimize the risk of overlooking relevant data, we retrieve the top five most similar tables.color=green]Is “asymmetric semantic similarity” a research field, a technique, or a particular model (maybe in GPT suite or library)? Need to be clear on this and provide link/reference. — a type of model determined by how it is trained, i.e., short query vs long document
Querying Relevant Tables.
Querying the relevant tables presents challenges, as the system must recognize how values in the claim are represented in the database. color=cyan]We said “presents challenges” quite many times in the system. You might want to consider add some variety to the wording. In cases where direct mappings are available (e.g., in the Vote frame case study), we map FEs to specific tables. However, in the OECD case study, where no direct mapping exists between FEs and table columns, the system must predict this mapping. To handle these cases, we employ LLM-based code generation, allowing for the execution of multiple simple queries using programmed logic, rather than relying on complex text-to-SQL generation.color=cyan]Should we include some text-to-SQL references?
@@
We represent the table schema as SQL code, following best practices from previous work Gao et al. (2023), which studied different methods of representing SQL tables for text-to-SQL generation.color=green]What is Gao et al. 2023 about? on the same problem or something else. YOu might want to make this clear. Alongside the schema, we include representative example values from the database for each column. color=green]“Columns …. include all values” doesn’t seem correct. — referring to how they are encoded into text Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values. For columns with a moderate number of unique values, we use a semantic similarity model to select the top ten most relevant values based on the claim. Given the large number of cells generated by this process, we use an LLM to refine the query and filter out irrelevant columns, focusing on those most critical for fact-checking the claim. The prompt for this LLM can be found in Appendix D.3. color=green]Is this semantic similarity model the same as the aymmetric one mentioned in the last section? If yes, make it clear; if not, need reference/link, based on the same concern mentioned earlier. — each model is different because they perform differently on the data, the selection is discussed in the experiments color=green]Are we missing some details here about how we used LLM? Are there prompts?
Fact Verification
@@
In the final step, we used an LLM-based fact verification model. We represent the extracted evidence as a list of tab-separated values based on the data retrieved in the previous step. The model is instructed using the prompt in Appendix D.4 and outputs one of five verdicts: false, mostly false, half-true, mostly true, or true.color=cyan]If time allows, explain the connection to PolitiFact’s Truth-o-meter.
5 Experiments
@@ 5.1 Datasets
Arslan et al. (2020) introduced 11 manually defined semantic frames to extend the long-running Berkeley FrameNet project Baker et al. (1998). Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in this dataset along with the newly-defined frames. These frames can be found in Appendix 6.
PolitiFact Fact-checks
For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles collected as of April 2022. Each article contains a claim, a detailed fact-check, a verdict, and a list of sources. To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we identified 79 fact-checks that cite congress.gov and evoke the Vote frame. For each voting claim, we manually verified the bills referenced within the fact-check are related to the claim to ensure the accuracy of our dataset.
color=cyan]PolitiFact fact-checks and PolitiFact fact-check articles. Mention very close to each other. color=green]the claim cites congress.gov? color=green]“bill cited was referenced by the claim”. cited where? also the claim would reference a bill?
Additionally, we collected 68 articles that cite oecd.org for our analysis of OECD-related claims. Each OECD claim was manually verified and mapped to the statistics which can be used to fact-check it. These two sets of claim-evidence pairs—the voting-related claims and the OECD-related claims—serve as evaluation sets for our case studies. color=green]these articles cite the website? color=green]“To collect … we collect” color=green]Which two sets?
@@ 5.2 Frame-Semantic Parsing
For frame-semantic parsing, we combine the frame identification model developed by Devasier et al. (2024) with a frame element identification model based on AGED Zheng et al. (2023). By combining these state-of-the-art approaches, we unify frame and frame element identification into a single model. This model is then trained only on the fact checking frames introduced in Arslan et al. (2020).
To evaluate the performance of our model for the case studies, we compare it with a GPT-4o-mini model using OpenAI’s structured generation [^8] At the time of submission, GPT-4o-mini is the only model which supports structured generation.. We evaluate each model using accuracy on frame identification and exact match accuracy on FE identification in Table 1. We define a separate prompt for each case study in Appendix D.1.
5.3 Claim Coverage
One concern which may arise is regarding the applicability of our system for assisting in the fact checking process. To understand this, we aim to identify the coverage of a frame-based approach towards automatically fact-checking claims. Due to the lack of annotations for all of the fact-checking frames in Arslan et al. (2020), we utilize a zero-shot GPT-4o-mini model to identify the frames evoked in the claim of each PolitiFact fact-check (Table 2). Claims which do not evoke a frame studied in this work are categorize as none. The prompt used for this model can be found in Appendix D.1.
@@ 5.3 Claim Coverage
text-embedding-3-large Vote 0.032
stella_en_1.5B_v5 Vote 0.144
distilbert-tas-b Vote 0.165
@@
distilbert-tas-b Full claim Vote 0.143
distilbert-tas-b Issue FE Vote 0.165
Max Possible - Vote 0.568
Table 4: Performance of different methods of representing the query in the OECD case study’s table matching. OECD and Vote claims use Recall@5/10, respectively.
5.4 Evidence Retrieval
We study the capability of our system to retrieve the relevant information which can be used to fact-check the claim. To do this, we found the percentage of bills and statistics available in our database for voting and OECD-related claims, respectively. This serves as a performance ceiling on our collected claims as our system would be unable to fact-check claims which do not have available data. These values are found in Table 4 (Max Possible).
For semantic similarity, we compare commonly used embedding models (Table 3), namely RoBERTa Liu et al. (2019) and DistilBERT-TAS-B Hofstätter et al. (2021). Additionally, we compare with OpenAI’s text-embedding-3-large model and provide a baseline BM25 approach. We also experiment with different approaches of representing the query used to find relevant data. In Table 4, we compare the performance of using the entire claim with only the relevant FE. To evaluate the performance of these queries, we calculate the recall@K, where K=5 for OECD claims and K=10 for Vote claims. The relevant FEs used for each frame can be found in Appendix A.
@@ 5.4 Evidence Retrieval
Our w/ Irrelevant Vote 0.076
Our w/o Irrelevant Vote 0.207
Our w/ Irrelevant ± Vote 0.253
GPT-4o Naive ± Vote 0.324
Our w/o Irrelevant ± Vote 0.690
@@
Our w/ Irrelevant OECD 0.214
Our w/o Irrelevant OECD 0.429
GPT-4o Naive ± OECD 0.537
Our w/ Irrelevant ± OECD 0.607
Our w/o Irrelevant ± OECD 0.821
Table 5: Fact verification performance on different case studies. ± indicates the off-by-one performance. Bolded values indicate the best-performing settings overall and underlined values are only considering exact matches.
@@ 5.5 Fact Verification
Finally, we study the ability of our fact verification system to fact-check claims using structured data by comparing the verdicts presented in each PolitiFact fact-check to the model’s predicted verdict. We have replaced PolitiFact’s “pants on fire” verdicts with a false verdict. To determine whether the LLM is relying on internal knowledge or if it is using the provided data, we evaluate the performance of GPT-4o without any data. We measure accuracy as the primary evaluation metric and also calculate off-by-one accuracy, where, for example, a prediction of “mostly false” is considered correct if the label is “false.” The results of these experiments are summarized in Table 5.
6 Results
6.1 Frame-Semantic Parsing
We evaluated the frame-semantic parsing model on fact-checking frames introduced by Arslan et al. (2020), comparing it against GPT-4o-mini due to its structured output capability. Table 1 shows the performance on both the Vote and OECD frames. We found that the GPT-4o-mini model tends to over-predict the number of frames in a sentence. When we only use the first predicted frame, the performance is increased significantly. We saw similar cases with frame element predictions where GPT-4o-mini tends to predict more frame elements than actually exist in the input. We include examples of each of these cases in Appendix C.2.
6.2 PolitiFact Survey
Similar to Section 6.1, we found that GPT-4o-mini tends to predict more frames than there may actually be. We believe this is likely due to the annotations from Arslan et al. (2020) being sampled from PolitiFact and likely having a similar structure. We found the annotated samples evoked 1.1 frames on average while GPT-4o-mini predicts an average of 2.1 frames. Another possible explanation is that the model is able to identify frames for lexical units– words which evoke frames–which are not defined.
One indication of this is on the Vote frame, which only has a lexical unit defined for vote.v; however, it is possible to evoke the Vote frame using other words. For example, the statement “I passed a bill” implies that a vote of affirmation was cast on that bill. We found only 1,198 (5.7%) PolitiFact fact-checks which mention “vote” while the model predicts that almost 3 times as many evoke the Vote frame. While we found many claims which evoke a Vote frame without an explicit mention of the word “vote”, further analysis is needed to get a more accurate understanding of the true distribution of frames evoked by factual claims.
6.3 Evidence Retrieval
We found that embedding models’ performance is task-specific and there is no single-best embedding model for every task. For example, RoBERTa performed the best for finding the right OECD table while DistilBERT-TAS-B was better for matching claims to their corresponding bills. We also found that, in general, using the claim-specific frame element performs better than using the entire claim as a query, as shown in Table 4.
For voting-related claims, we found that identifying the bill referenced in a particular claim is very difficult. Our best performing model, DistilBERT-TAS-B, was only able to obtain a Recall@10 of 0.165. For OECD-related claims, identifying the table to use to fact-check the given claim was much easier, with the best model, RoBERTa, achieving a recall@5 of 0.726. This is likely because the similarity between the relevant FE and the table names were better-represented by semantic similarity models. This indicates a potential area of significant improvement in the methodology and representation of bill texts according to their topics and implications, in addition to the text itself.
6.4 Fact Verification
On OECD claims, 62% of the generated queries were able to retrieve relevant data while only 36% of the voting claims were able to retrieve relevant bills. These values put an upper bound on the capabilities of our fact verification model as without evidence, the model will not be able to provide useful fact-checks. In general, we found that language models tend not to predict a claim to be true based on the given data, yet they have no issue with predicting a claim to be false, and often do so. We believe this is analogous to proofs, it is quite easy to prove by contradiction, if you find a single vote or country statistic which contradicts the claim, while it is much harder to assert that a claim is true without looking at all of the data.
We found that on Vote claims, a naive GPT-4o correctly predicted the truthfulness of the factual claim with an accuracy of 0.286; however, when voting data is added, the performance drops to 0.076. This suggests that, while there may be some internal knowledge on certain people and bills, it is unlikely that the model is relying on it when given specific voting records. This also indicates that the prompt given to the model is sufficient in preventing the LLM from relying on internal knowledge.
When excluding claims for which data was not found, we observed absolute performance increases of +13.1% and +21.5% for Vote and OECD claims, respectively. We also found that, while the exact-match accuracy of the system is low, there is a strong correlation between predictions and labels, as indicated by the off-by-one accuracy in Table 5. We include example outputs for the explanations of these models in Appendix C.1.
@@ 7 Conclusion
This research contributes to the evolving field of automatic fact-checking by proposing a novel paradigm that incorporates frame semantics to enhance the evidence retrieval process and enable fine-grained control of downstream tasks. By developing and presenting a pilot dataset specifically annotated for real-world claims, this work helps to address the underexplored challenge of fact-checking using massive, structured data sources.
Our findings demonstrate the promise of frame-semantic parsing in improving evidence retrieval for automatic fact-checking. In both case studies, we observed that using frame elements extracted from claims instead of entire claims led to substantial improvements in recall metrics, highlighting the utility of frame-based understanding in identifying relevant evidence. These results suggest that semantic frames not only offer a structured and explainable approach to parsing claims but also guide retrieval models more effectively toward the correct information within large datasets.
Looking ahead, the integration of frame semantics into fact-checking systems holds promise for further advancements in both accuracy and interpretability. Future work could explore broader applications of this approach, including expanding the dataset, refining the models, and incorporating additional semantic frames to cover more claims.
@@
Despite the advancements made in this study, our system’s limitations should be acknowledged. While this study focuses primarily on improving evidence retrieval and frame-semantic parsing, it does not comprehensively evaluate different fact verification approaches. Fact verification is critical for automatic fact-checking, and a deeper investigation into various verification strategies, including fine-tuned models and LLM-based approaches, could further enhance system performance.
@@
Appendix A Studied Frames
Table 6 provides a comprehensive list of frames studied in this work along with the frame elements used to predict table mappings.

### v2

@@ 4.2 OECD Statistics
The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics which serve as a strong focal point to explore a wider range of semantic frames. For the OECD case study, we collected all of the data tables available on the OECD Data Explorer.  [^7] https://data-explorer.oecd.org/ We constructed an SQL database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totaling over 4.1 billion cells.
@@
To identify the relevant table(s) to fact-check a given claim, we use a RoBERTa-based semantic similarity model Liu et al. (2019). To represent the table, we encode the table’s name and description as text. To minimize the likelihood of missing relevant tables, we retrieve the five tables most similar to the extracted frame element from the claim.
Querying retrieved tables.
Querying the retrieved tables requires understanding of the tables, columns, and the values within them. To do this, we encode each table’s name, all of its columns, and several representative sample values from each column into text. We then utilize GPT-4o to understand the requirements for verifying a claim and generate Python code to query the retrieved table accordingly. We choose this approach, as opposed to standard text-to-SQL approaches because it allows for the execution of multiple simple queries using programmed logic, rather than relying on text-to-SQL generation which can result in very complex queries for simple claims.
@@
We represent the table schema as SQL code, following best practices from previous work Gao et al. (2024). Alongside the schema, we include representative example values from the database for each column. Our implementation is similar to Nahid and Rafiei (2024); however, we address critical limitations to support the high-volume OECD data. Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values. For other columns we use the same RoBERTa-based semantic similarity model to select the top ten most relevant values based on the claim. These queries often result in a large number of cells, so we instruct GPT-4o to refine the query by filtering out irrelevant columns, focusing on those most critical for fact-checking the claim. The prompts for this process can be found in Appendix E.3.
Fact verification.
@@
In the final step, we used an LLM-based fact verification model. We represent the extracted evidence as a list of tab-separated values using the data retrieved in the previous step. The model is instructed using the prompt in Appendix E.4 and outputs one of five verdicts: “false,” “mostly false,” “half-true,” “mostly true,” or “true,” based on PolitiFact’s Truth-o-Meter.  [^8] https://www.politifact.com/article/2018/feb/12/principles-truth-o-meter-politifacts-methodology-i/
5 Experiments and Results
@@ 5.1 Datasets
Arslan et al. (2020) introduced 11 manually defined semantic frames to extend the long-running Berkeley FrameNet project Baker et al. (1998). Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in their dataset along with the newly-defined frames. These frames can be found in Table 6 in the Appendix.
PolitiFact fact-checks.
For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles (i.e., fact-checks) collected as of April 2022. Each article contains a claim, a detailed fact-checking analysis, a verdict, and a list of sources. To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we manually identified 79 fact-checks that cite congress.gov in the sources and evoke the Vote frame. For each voting claim, we manually verified the bills referenced within the fact-check are related to the claim to ensure the accuracy of our dataset. Similarly, we collected 68 fact-checks that cite oecd.org for our analysis of OECD-related claims. Each OECD claim was manually verified and mapped to the database table which can be used to fact-check it. Additional details are provided in Appendix A.
@@ 5.2 Frame-Semantic Parsing
For frame-semantic parsing, we combined the frame identification model developed by Devasier et al. (2024) with a frame element identification model based on AGED Zheng et al. (2023). By combining these state-of-the-art approaches, we unified frame and frame element identification into a single RoBERTa-based model (additional details in Appendix C.4). We fine-tuned two of these frame-semantic parsing models, one for each case study, using their respective frames from the training data in Section 5.1. The model used by our system is then selected based on the candidate targets in the sentence (extracted using the approach in Devasier et al., 2024). To evaluate the performance of our model for the case studies, we compared it with a baseline GPT-4o-mini model using OpenAI’s structured generation. The prompt used for this model can be found in Appendix E.1. We evaluated each model using exact match accuracy on frame identification and argument identification. These results are presented in Table 1.
Our findings are consistent with a previous study indicating that generative LLMs tend to struggle with frame-semantic parsing Su et al. (2024). Our fine-tuned frame-semantic parsing model performed much better on frame and argument identification for both case studies. A key observation was GPT-4o-mini’s tendency to over-predict frame occurrence in sentences. Performance improved significantly when we constrained the analysis to only the first predicted frame (GPT-4o-mini*). Similar patterns emerged in frame element predictions, where GPT-4o-mini frequently predicted frame elements that do not exist in the frame definition or are not used in the input sentence. Examples of these are provided in Appendix D.2.
5.3 PolitiFact Survey
To assess the broader applicability of our frame-semantics-based fact-checking system, we conducted a comprehensive coverage analysis. Because the annotations available for the fact-checking frames are limited, we employed a zero-shot GPT-4o-mini model (detailed in Appendix E.1) to identify frames evoked in the PolitiFact fact-check corpus. Claims that did not trigger any of our studied frames were classified as “None.” The distribution of frames is presented in Table 2.
Our findings show that GPT-4o-mini may predict more frames than there actually are. This is because the annotated dataset evokes an average of 1.1 frames per claim while GPT-4o-mini predicted an average of 2.1 frames per claim. We believe this is because GPT-4o-mini predicts frames for lexical units–words which evoke frames–that are not defined by Arslan et al. (2020).color=yellow]This last sentence appears to suggest GPT can find frames there were missed in Fatma’s dataset. But why did you say it “predict more than there actually are”? — This isn’t really known 100% the cause could be that the model over-predicts in a bad way or it could be because the model recognizes frames for more LUs than are defined. It seems to be the latter based on what ive seen.
One indication of this can be seen with the Vote frame. While it explicitly defines only the lexical unit vote.v, the frame could be evoked through various other expressions. For instance, the statement “I passed a bill” implicitly suggests a vote of affirmation. Our analysis found that only 1,198 (5.7%) of PolitiFact fact-checks explicitly mention “vote,” whereas the GPT-4o-mini model predicted Vote frames at nearly three times that amount. Although we observed numerous claims that evoked the Vote frame without explicitly using the word “vote,” additional research is needed to determine the true distribution of frames in factual claims.
@@ 5.3 PolitiFact Survey
text-embedding-3-large Vote 0.032
BM25 Vote 0.048
roberta-base-v2 Vote 0.131
stella_en_1.5B_v5 Vote 0.144
distilbert-tas-b Vote 0.165
@@
5.4 Evidence Retrieval
Similarity model selection.
To retrieve relevant information for fact-checking, we first evaluated different semantic similarity models. Table 3 presents the embedding models we compared, including Stella Zhang et al. (2025), RoBERTa Liu et al. (2019), DistilBERT-TAS-B Hofstätter et al. (2021), OpenAI’s text-embedding-3, 999 https://platform.openai.com/docs/models and BM25. 101010 https://github.com/dorianbrown/rank_bm25 color=yellow]In Table 4, font size of different models are different; some R@K is italic, others not. Table 1 and Table 3/4 use different font styles.
Our analysis revealed that different models excel in specific verification contexts. RoBERTa demonstrated superior performance in identifying relevant OECD tables, while DistilBERT-TAS-B achieved the best results in matching claims to bills. These findings suggest that optimal fact-checking systems should employ task-specific embedding models rather than a one-size-fits-all solution.
@@ 5.4 Evidence Retrieval
distilbert-tas-b Full claim Vote 0.143
distilbert-tas-b Issue FE Vote 0.165
Max Possible - Vote 0.568
@@
Table 4: Performance of different query representations in each case study. Evaluations are performed on table matching (OECD) and bill matching (Vote), and use Recall@5/10, respectively.
Query representation.
We explored two different methods for constructing retrieval queries: using the complete claim text and using only the frame elements (FEs) extracted based on the evoked frame (detailed in Appendix B). We evaluated these query representations using the best-performing similarity models identified in Table 3.
Table 4 presents the recall@K metrics for each representation approach, with K=5 for OECD claims and K=10 for Vote claims. We also established baseline metrics for evidence availability (Max Possible) in our database. For voting-related claims, we calculated the percentage of referenced bills present in our database, while for OECD claims, we assessed the availability of relevant datasets.
Our results demonstrate that queries constructed from frame elements consistently outperformed full-claim queries across both domains. This finding aligns with research by Madabushi (2024) showing that frame-semantics-based search strategies enhance retrieval effectiveness.
The performance varied significantly between domains. For voting-related claims, even our best-performing model (DistilBERT-TAS-B) achieved a modest recall@10 of 0.165, reflecting the complexity of bill matching. In contrast, OECD claim verification was easier for the model, with RoBERTa achieving a recall@5 of 0.726, likely due to the more constrained search space. Congress member identification performed very well with an exact match accuracy of 97.5%.
Query generation.
The effectiveness of generated queries varied significantly between the two domains: 62% of OECD claim queries successfully retrieved relevant data, compared to only 36% for voting claims. These retrieval rates effectively establish upper bounds for our fact verification capabilities, as meaningful fact-checking requires successful evidence retrieval.
@@
Ours w/ Irrelevant Vote 0.076
Ours w/o Irrelevant Vote 0.207
@@
Ours w/ Irrelevant OECD 0.214
Ours w/o Irrelevant OECD 0.429
Table 5: End-to-end fact verification performance on the two case studies. The best-performing models are bolded.
@@ 5.5 Fact Verification
We evaluated our fact verification system by comparing its predicted verdicts against the ground truth verdicts from PolitiFact fact-checks (Table 5). For consistency in our analysis, we consolidated PolitiFact’s "pants on fire" rating with "false".
To assess whether our system genuinely relies on retrieved evidence rather than pre-existing knowledge, we established a baseline using GPT-4o without access to any external data (GPT-4o Naive). Our analysis encompassed claims across both voting records and OECD datasets, with performance measured against human-annotated verdicts.
The baseline GPT-4o Naive model performed poorly on both case studies, suggesting that the model’s internal knowledge alone is insufficient for accurate fact verification. When we excluded cases where our system failed to retrieve relevant evidence (w/o Irrelevant), we observed substantial improvements in accuracy: an increase of 13.1 percentage points for voting-related claims and 21.5 percentage points for OECD-related claims. This improvement demonstrates the critical role of successful evidence retrieval in fact verification accuracy. Detailed examples of the system’s reasoning and explanations can be found in Appendix D.1.
Our analysis revealed a bias in the model’s predictions toward classifying claims as false. This bias likely stems from the nature of contradiction-based verification where disproving a claim requires finding just one contradictory piece of evidence, while proving a claim often necessitates comprehensive evidence supporting all aspects of the assertion. This suggests potential areas for future improvement in balancing verification criteria.
@@ 6 Conclusion
In this work, we introduced a novel paradigm for task-oriented automatic fact-checking by leveraging frame semantics to enhance structured understanding and evidence retrieval. Through the development of a pilot dataset composed of real-world claims and two focused case studies, we demonstrated how frame-semantic parsing can guide the fact-checking process with structured data, improving retrieval performance and explainability. Our findings suggest that semantic frames offer a meaningful way to bridge the gap between complex claims and high-volume structured datasets.
The empirical results highlight the advantages of using frame elements for evidence retrieval, with improvements in recall across both voting-related and OECD-based fact-checking tasks. By aligning claims with relevant structured data, our approach not only enhances accuracy but also provides a more interpretable verification process. Furthermore, our survey of fact-checking frames reveals promising avenues for future research, particularly in expanding the coverage of high-impact frames and refining claim-evidence alignment techniques.
While our system is limited by the availability of trustworthy data, our findings highlight a critical research gap left by approaches that focus on smaller tabular data. This work showcases the use of high-volume structured databases for automatically verifying novel claims. Expanding beyond our pilot datasets to large-scale collections of claims and corresponding evidence from structured sources remains an important next step.
@@
Despite the advancements made in this study, our system’s limitations should be acknowledged. While this study focuses primarily on improving evidence retrieval, it does not comprehensively evaluate different fact verification approaches. Fact verification is critical for automatic fact-checking, and a deeper investigation into various verification strategies, including fine-tuned models and LLM-based approaches, could further enhance system performance.
@@
Appendix A Data Quality